For questions 1 consider the Maclaurin Series expansion of



x = + + + +1 x x2 x3 .....



e



2! 3!



1. The coefficient of the term containing x3 in the Maclaurin series expansion of e3x



is 3





1−e3x




1
Expert's answer
2021-11-24T17:39:36-0500

"e^x=1+x+\dfrac{x^2}{2!}+\dfrac{x^3}{3!}+..."

1. Then


"e^{3x}=1+3x+\dfrac{(3x)^2}{2!}+\dfrac{(3x)^3}{3!}+..."

"\dfrac{(3x)^3}{3!}=\dfrac{9}{2}x^3"

The coefficient of the term containing "x^3" in the Maclaurin series expansion of "e^{3x}" is "9\/2."


2.


"\lim\limits_{x\to0}\dfrac{1-e^{3x}}{x}"

"=\lim\limits_{x\to0}\dfrac{1-(1+3x+\dfrac{(3x)^2}{2!}+\dfrac{(3x)^3}{3!}+..)}{x}"


"=-\lim\limits_{x\to0}\dfrac{3x+\dfrac{(3x)^2}{2!}+\dfrac{(3x)^3}{3!}+..}{x}"

"=-\lim\limits_{x\to0}(3+\dfrac{9x}{2!}+\dfrac{27x^2}{3!}+...)=-3"